OpenStudy (anonymous):

Determine the domain and range of each function. Write in interval notation. f(x)=x^2-5

OpenStudy (solomonzelman):

Well, the domain (the x value you plug into the function, for x) can be any number. It can be any positive or negative real number. Since there are no restrictions. Domain (-∞,∞) For the RANGE: You can see that if x = 0 then you can get a range of -5 (this the y-intercept) You won't get a smaller range than `-5` , see why ? RANGE [-5,∞)

OpenStudy (solomonzelman):

Does this make sense ?

OpenStudy (anonymous):

yes. thank you. so for f(x)=-\[\sqrt{x-3}\] domain: [-3,\[\infty\]) range: (-\[\infty\]]

OpenStudy (anonymous):

sorry still new to this whole setup for the equation typing

OpenStudy (solomonzelman):

You can type an infinity symbol, do you want to know how ?

OpenStudy (solomonzelman):

I am on the keyboard and anywhere..... 1) Click and hold ALT 2) click the number code ` 2 3 6 ` (using the numbers that are on the right of the keyboard, not the ones below F1, F2, F3, etc., ) 3) release the ALT (it should show ∞ )

OpenStudy (anonymous):

∞ got it. thanks

OpenStudy (solomonzelman):

Okay, but for \(\normalsize\color{blue}{ f(x)=-\sqrt{x-3}}\) The domain would be \(\normalsize\color{blue}{ [3,∞)}\) because if it is less than 3 you get an imaginary number. The range will be \(\normalsize\color{blue}{ [0,∞)}\) because the smallest value you can get for `f(x)` is a 0 (at x=3) .